OpenStudy (anonymous):

Roberto is going to visit his family in Mexico. He is bringing $500 US dollars. On the way over, he can exchange his US Dollars for Pesos. The exchange rate is $0.07 USD = 1 Peso. How much money will he have to spend in Mexico, according to their monetary system? A. 35 Pesos B. 700 Pesos C. $0.0004 Pesos D. $7142.86 Pesos I got B as my answer, would that be correct?

Directrix (directrix):

.07/1 = 500/x where x is the number of pesos .07x = 1*500 .07x = 500 x = 500/.07 x = ? @katlin95

Directrix (directrix):

> I got B as my answer, would that be correct? No. Try again. @katlin95